OpenStudy (anonymous):

Find dx/dy in terms of y when y(x)=(e^x)-(e^-x) I know dy/dx=(e^x)+(e^-x) I'm trying to get x in terms of y.

OpenStudy (loser66):

he asks for \(\dfrac{dx}{dy}\)

OpenStudy (loser66):

to me, I convert \(sinh x = \dfrac{e^x-e^{-x}}{2}\) \(\rightarrow \) 2 sinh x = \(e^x -e^{-x}\) =y

OpenStudy (loser66):

from then, I have y/2 = sinh x \(\rightarrow\) x = 1/2 arcsinh (y) then, take derivative , you have dx/dy = \(\dfrac{1}{2(\sqrt{1-y^2)}}\)

OpenStudy (anonymous):

May I ask, where did you get sinhx from?

OpenStudy (loser66):

that is the form of it, kind of formula. That question is equivalent to "why 1+1=2"

OpenStudy (loser66):

look at this http://tutorial.math.lamar.edu/Classes/CalcI/DiffHyperTrigFcns.aspx

OpenStudy (anonymous):

I see, so I have to dx/dy : 1/2 arcsinh y.

OpenStudy (loser66):

you have formula for this, too. the problem is just convert, and that's all.

OpenStudy (anonymous):

Can you show me how this is done? or the relevant formula. Thanks.

OpenStudy (loser66):

look at this http://math.info/Calculus/Derivatives_Hyp_InvHyp/

OpenStudy (anonymous):

The answer is \[\frac{ 1 }{ \sqrt{y ^{2}+4} }\] It is also possible to do this by taking \[y ^{2} and (\frac{ dy }{ dy })^2 \] and comparing the two results.

OpenStudy (loser66):

don't get what you mean